OpenStudy (anonymous):

If I varies inversely as r+2, and I=9 when r=3, find r when I=7. Help please?

hero (hero):

The usual formula for inverse variation is xy = k In this case, l = x y = r + 2 so l(r+2) = k which means 9(3+2) = k 9(5) = k 45 = k Now use the same formula in terms of l and r to find the new value for r given that l = 7.